# One Queue Is All You Need: Resolving Head-of-Line Blocking in Large Language Model Serving

### v1

@@
Abstract
Large language models (LLMs) have become an increasingly important workload for cloud providers catering to both enterprise and consumer applications. LLM inference requests from these applications have end-to-end latency SLOs that must be adhered to in production settings. However, existing LLM serving systems focus on optimization objectives such as request serving throughput or request execution latency rather than the end-to-end latency SLOs. Achieving end-to-end SLOs for latency-sensitive requests is challenging due to head-of-line (HOL) blocking in the request queue, which results from bursty arrival rates and insufficient resources.
To address the above challenge, we propose QLM, a multi-model queue management framework for LLM serving. QLM uses stochastic programming to orchestrate the actions of multiple LLM Serving Operations (LSOs) to reduce HOL blocking and maximize SLO attainment. Specifically, QLM uses the following LSOs: model swapping, request eviction, GPU-CPU state swapping, load balancing, and warm model start. Evaluation on heterogeneous GPU devices and models with real-world LLM serving dataset shows that QLM improves SLO attainment by 40–90% and throughput by 20–400% while maintaining or improving device utilization compared to other state-of-the-art LLM serving systems.
1 Introduction
Motivation. The emergence of large language models (LLMs) such as OpenAI GPT-4 and Google Gemini has enabled a wide range of novel AI applications [58, 3, 56], including chatbots and coding assistants. Consequently, serving LLMs has become an increasingly critical workload, catering to both enterprise and consumer applications with service-level objectives (SLOs) on end-to-end latency [20, 37, 53]. However, existing LLM-serving systems [21, 57, 48, 49, 38, 25, 45, 55] focus on optimization objectives such as request serving throughput, device utilization, or request execution latency rather than addressing SLOs on the end-to-end latency, which encompasses both request execution time and waiting time. Optimizing end-to-end latency SLOs is critical as it is the prime metric valued by the customers using LLM inferencing services [31, 30].
Figure 1: QLM nearly eliminates HOL blocking time in both single and multi-model serving scenarios and enables SLO attainment for batch and latency-sensitive applications.
When maximizing end-to-end latency SLO attainment, minimizing request waiting time is just as important as execution time due to the presence of request queues. Request queues fill up when the serving throughput is unable to match the high and bursty arrival rates of the incoming requests [53]. The issue of increasing request queue sizes is further exacerbated by device shortage and the need to serve multiple fine-tuned heterogeneous models with varying sizes [24], resulting in high cold start time [12] and low device utilization [62]. Latency-sensitive requests with strict SLOs may wait longer in the queue because of head-of-line (HOL) blocking, leading to potential SLO violations. Figure 1 (left) shows that requests can suffer from HOL blocking in both single and multi-model serving using production LLM-serving traces [36] (labeled as “vLLM”, a state-of-the-art LLM serving system).
Our Work. Fundamentally, HOL blocking can be alleviated by reordering the requests in the waiting queue and assigning requests to appropriate devices. In the LLM serving context, this assignment and reordering can be mapped to LLM serving operations (LSOs) such as request pulling, request eviction, load balancing, GPU-CPU state swap, and model warm start. To orchestrate these LSOs and maximize SLO attainment, we introduce QLM, a multi-model queue management system. Figure 1 (right) shows that QLM maximizes SLO attainment for both latency-sensitive and batch jobs in a multi-model serving workload setup. QLM leverages two major ideas in its queue management framework:
  1. (a)
    Virtual Queue Abstraction: Previous work tends to optimize LSOs separately and in an ad-hoc manner [54, 15, 46]. An LLM serving framework that systematically investigates the trade-offs amongst multiple LSOs for SLO attainment is missing to date. To bridge this gap, QLM introduces the concept of virtual queues that represent the dynamic order and assignment of requests to be served, which provides a necessary abstraction for optimizing LSO actions. Moreover, selecting optimal LSOs per request is an NP-hard decision problem that will not meet the acceptable overhead of executing requests in production (e.g., < 10 ms per request). To reduce the complexity of the decision problem, QLM uses request groups, where requests that share the same performance characteristics are grouped together, and decisions are taken per request group instead of per request.
  2. (b)
    Plan Generator and Request Completion Time (RCT) Estimator: When making decisions on request group ordering in the virtual queues, the two key metrics that need to be estimated are the request waiting time and execution time. However, estimating these metrics is challenging because the execution time for a request depends on its output token length, which is unknown apriori before executing the request [21] (i.e., stochastic in nature). QLM models this stochastic nature and estimates the waiting and execution time distributions using a Request Completion Time (RCT) estimator (described in Section 6). Finally, QLM uses a Plan Generator (described in Section 7) that leverages the completion time distribution of request groups to create an optimal ordering and assignment of request groups onto the virtual queues to maximize SLO attainment. Because the key variables in the optimization, output token length and completion times are stochastic variables, the plan generator is based on a stochastic programming solver. In comparison, other ML serving systems such as Clockwork [16] and SHEPHERD [60] use variations of linear programming solvers because the request completion time and waiting time are deterministic for traditional ML models (e.g., ResNet).
    In summary, QLM enables the translation of end-to-end per-request latency SLOs into backend LSO actions using a combination of virtual queues, stochastic modeling of queue dynamics (in RCT estimator), and stochastic programming (in plan generator).
Results. We demonstrate QLM on our internal production-grade version of vLLM [21] as the backend LLM-serving system. QLM supports the following five basic LSOs (see section 5 for details): (1) Request Pullingfrom the global waiting queue into the running batch in the GPU, (2) Request Evictionfrom the running batch back into the waiting queue, (3) GPU-CPU Swappingfor the internal LLM state, (4) Model Warm Startfrom CPU memory instead of disk, and (5) Load Balancingacross multiple LLM model instances.
@@ 1 Introduction
We evaluate QLM on three popular LLMs of varying sizes (i.e., Mistral-7B [19], Vicuna-13B [5], and Llama-70B [50]) on heterogeneous GPU clusters with NVIDIA A10 and A100 GPUs. We adopt workloads from a real-world LLM dataset: ShareGPT [43] using setups derived from our production requirements. Our experiments demonstrate the following major improvements with QLM:
@@
    SLO Attainment: QLM achieves 40–90% higher SLO attainment compared to the vanilla vLLM serving system and 50–90% higher SLO attainment compared to traditional ML serving systems like SHEPHERD,
@@
    Request Throughput: QLM improves the request throughput in a multi-model serving system by 400% on average and in a single-model serving system by 20% on average compared to other LLM serving systems, and
@@
    LSO Ablation Study: QLM demonstrates that all LSOs contribute to SLO attainment and throughput improvement. Notably, we find that model warm start improves throughput by 300% in multi-model serving, and request eviction improves SLO attainment by 80% in single-model serving.
2 Background
2.1 LLM Inference
@@
Continuous Batching. During LLM inference, the decoding stage is memory-bound, as loading model weights from memory takes longer than computation. Therefore, state-of-the-art LLM serving systems like vLLM [21], Orca [57], Tensor-RT [48] and TGI [49] employ continuous batching with iterative scheduling to enable dynamic addition of requests to a batch as soon as others have finished generation.
@@
PagedAttention. Static allocation of the KV cache can result in significant memory waste as the KV cache grows dynamically during the decoding stage. PagedAttention [21] introduces the idea of managing the KV cache, like OS memory, via pages and enabling dynamic allocation. Such dynamic allocation prevents fragmentation and enables nearly 100% utilization of GPU memory and furthers throughput improvement when combined with continuous batching [21].
Figure 2: Interaction between LLM applications and a multi-model LLM serving system.
2.2 LLM Serving Systems
@@
Figure 2 shows the typical workflow of an LLM serving system. User-facing applications such as chatbots, log processing, and recommenders have specific end-to-end latency SLO requirements [37, 20, 39, 4]. When interacting with LLMs, each application generates requests that consist of the input prompts and associated metadata (e.g., model type and SLO value) to the LLM serving system. For example, chatbots require requests to complete by a deadline (e.g., p99 completion time <10s [14]), and batch jobs like log processing have a more relaxed SLO. Note that end-to-end SLOs do not specify per-token metrics such as time per output token (TPOT) or time to first token (TTFT) [10] as they cannot be directly translated to the application’s latency requirements.
These requests are then served by an LLM serving system such as vLLM [21], TGI [49], and Nvidia Triton [51] that implement state-of-the-art inference optimizations such as continuous batching and paged attention. However, even though these systems are optimized for high throughput serving, served requests can suffer from head-of-line (HOL) blocking [52, 38] when the request arrival rate exceeds the LLM serving system throughput [53]. In such cases, request queues would form, and requests with strict SLOs at the end of the queue would have to wait for earlier requests to complete, resulting in potential SLO violations.
Managing these request queues is further complicated by the growth of specialized, fine-tuned models [18, 44] for various LLM-augmented applications. For example, Code Llama [42] is fine-tuned for coding assistance, and Llama-chat [50] is fine-tuned for chatbots. As maintaining a standalone LLM serving system for each of these models can be expensive, they often have to be multiplexed together sharing the same serving system [6, 25, 22, 44]. However, serving multiple heterogeneous models with variable sizes results in high cold start time (that is needed for loading the model to the device memory) and low device utilization.
LLM serving systems have to manage these requests and models with limited devices through various backend LLM serving operations (LSOs) as defined in definition 2.4 such as request eviction [54], load balancing [15], and GPU-CPU state swap [46], in order to meet the SLO requirements. Managing these LSOs can be complex because of (1) their inter-dependency with each other, and (2) the auto-regressive nature of LLMs where the length of output tokens is highly variable.
2.3 QLM Definitions
@@
SLO: Each request arrives with a service-level objective (SLO) value that enforces the end-to-end time required to complete the request (request waiting time and execution time) [30]. Note that SLOs in QLM are defined per query and not per token (e.g., TPOT or TTFT [10]). End-to-end request completion time is the prime metric valued by the customers using LLM inferencing services [31, 30] as it can be directly converted into downstream application SLOs.
@@
LLM Serving Instance: An LLM serving system [^1] The difference between an LLM serving system and an LLM serving instance is similar to JAVA classes and objects. For example, vLLM is an LLM serving system while vLLM with a loaded model like Llama 70B is an instance of the LLM serving system. is capable of hosting LLM models by providing the necessary infrastructure and resources to load the models into memory and respond to requests. QLM is compatible with existing LLM serving systems such as vLLM [21] and TGI [49]. An LLM serving instance is composed of the LLM serving system and an LLM model that is being served.
Definition 2.4.
LLM Serving Operations (LSOs): In each LLM serving instance, the models and requests in the queue are managed by the LLM serving operations (LSOs). Examples of LSOs include request eviction [54], load balancing [15], GPU-CPU state swap [46]. Each LSO can trigger an action that either depends on or alters the state of the request queue (see section 5 for details).
2.4 Motivation and Characterization
Figure 3: Requests have predictable waiting times in a continuous batching system.
Figure 4: Choice between GPU-CPU state swapping and request eviction depends on the model setup.
Figure 5: Model swapping and request pulling can jointly decrease queue drain time.
To meet end-to-end application SLOs, it is critical to understand (1) the impact of LLM autoregressive patterns on the request completion time, (2) the complex interrelationship among individual LSOs, and (3) how end-to-end SLOs can be translated to actions for backend LSOs.
@@ 2.4 Motivation and Characterization
We characterize the performance of a state-of-the-art LLM serving system, vLLM [21], augmented with various LSOs to motivate the design of QLM. We use ShareGPT [43] and Azure LLM serving traces [36] from production environments. We present three key insights below.
Insight #1: Request waiting times can be accurately estimated with analytical methods. While individual request completion times in LLM inference can vary significantly, the average waiting time for a request in the queue is predictable. The waiting time can be estimated by dividing the total number of output tokens for requests ahead in the queue with the token generation throughput. Both average number of output tokens and throughput can be estimated by profiling the workload over time. We empirically find that these statistics do not change significantly for a workload and hence can be assumed constant. We validate the above waiting time estimation formula using real-world traces [43, 36]. Figure 5 illustrates this linear relationship between waiting time and queue position when serving requests for three varying-sized LLMs on NVIDIA A100 GPUs. Additionally, we find that the estimator is highly accurate with a coefficient of determination ($R^{2}$) of 0.99 (out of 1.0). In Section 6, we extend this estimator further to support request completion times estimation.
Insight #2: Selecting the optimal LSO depends on the model and hardware setup. When multiple LSOs are enabled for LLM inference, there can be conflicts between the goals of two LSOs. For example, when a request is blocked in the queue due to out-of-capacity on the GPU device serving the running batch, two LSOs can be used to allow immediate execution: request eviction and GPU-CPU state swapping. This choice primarily depends upon the trade-off between available CPU memory and the model/hardware-dependent computation cost of swapping vs. eviction. If additional memory consumption per KV cache token is lower relative to the KV recomputation time, then GPU-CPU state swapping would become preferable (and vice-versa).
To demonstrate this trade-off, we perform experiments with varying-sized models on NVIDIA A100 GPUs. Figure 5 shows the request throughput and CPU memory usage across these experiments. For models such as Mistral-7B, swapping increases request throughput by $\sim$20with a relatively small CPU memory overhead of 20 GB. On the other hand, for models such as Vicuna-13B, swapping increases CPU memory requirements by 8.4× (from 26 GB to 220 GB) with negligible throughput benefit. Therefore, the optimal LSO selection in this example would be to prefer eviction for Vicuna-13B and state swapping for Mistral-7B.
@@
Insight #3: Multiple LSOs can work together to attain SLOs while improving serving throughput. While some LSOs can have conflicting actuation actions (as shown in Insight #2), others, such as model swapping and request pulling, can work together to attain end-to-end latency SLOs while improving the throughput of the LLM serving system. Consider the case where each LSO is being optimized independently. The optimal request pulling strategy is to use the Earliest Deadline First (EDF) scheduling to achieve the maximum number of requests that meet their SLOs. However, this assumes that the model swapping cost is negligible. Frequent model swaps can happen (similar to thrashing) if multiple models are served to time share the same GPU devices, leading to SLO violations due to longer completion times to drain the queue and a drop in throughput. For example, consider the case illustrated in Figure 5. Requests with varying SLOs arrive in the queue, and they are placed by an EDF policy, causing multiple model swaps and substantially higher time to drain the entire request queue. Specifically, we find that across models and GPUs, the time required to serve all requests in the queue (i.e., the queue drain time) is substantially higher for the EDF policy compared to an Oracle policy that groups requests from the same model together to prevent the overhead of repetitive model swaps.
3 QLM Design Overview
Figure 6: Overview of QLM.
QLM aims to maximize the end-to-end latency SLOs by providing a translation layer between the SLOs and the actions for the LLM Serving Operations (LSOs).
3.1 Lifecycle of a Request in QLM

### v2

# Queue Management for SLO-Oriented Large Language Model Serving

@@
Abstract.
Large language model (LLM) serving is becoming an increasingly critical workload for cloud providers. Existing LLM serving systems focus on interactive requests, such as chatbots and coding assistants, with tight latency SLO requirements. However, when such systems execute batch requests that have relaxed SLOs along with interactive requests, it leads to poor multiplexing and inefficient resource utilization. To address these challenges, we propose QLM, a queue management system for LLM serving. QLM maintains batch and interactive requests across different models and SLOs in a request queue. Optimal ordering of the request queue is critical to maintain SLOs while ensuring high resource utilization. To generate this optimal ordering, QLM uses a Request Waiting Time (RWT) Estimator that estimates the waiting times for requests in the request queue. These estimates are used by a global scheduler to orchestrate LLM Serving Operations (LSOs) such as request pulling, request eviction, load balancing, and model swapping. Evaluation on heterogeneous GPU devices and models with real-world LLM serving dataset shows that QLM improves SLO attainment by 40–90% and throughput by 20–400% while maintaining or improving device utilization compared to other state-of-the-art LLM serving systems. QLM’s evaluation is based on the production requirements of a cloud provider. QLM is publicly available at https://www.github.com/QLM-project/QLM.
large language models, machine learning inference, queuing
journalyear: 2024copyright: rightsretainedconference: ACM Symposium on Cloud Computing; November 20–22,2024; Redmond, WA, USAbooktitle: ACM Symposium on Cloud Computing (SoCC ’24), November 20–22, 2024, Redmond, WA, USAdoi: 10.1145/3698038.3698523isbn: 979-8-4007-1286-9/24/11ccs: Computing methodologies Machine learningccs: Computing methodologies Distributed algorithmsccs: Computer systems organization Dependable and fault-tolerant systems and networks
1. Introduction
Motivation. Large language models (LLMs) such as OpenAI GPT-4 and Google Gemini have enabled novel capabilities in a wide range of AI applications (Zaharia et al., 2024; Bommasani et al., 2021; Wu et al., 2023a) such as chatbots and coding assistants. These base models such as GPT4 are further fine-tuned to support specialized tasks such as copywriting, financial planning, etc. (OpenAI, 2024). Consequently, serving multiple models for enterprise and consumer applications with latency-oriented service-level objectives (SLOs) has become increasingly critical (Jyothi et al., 2016; Qiu et al., 2020; Wang et al., 2024).
Previous work in this area (Kwon et al., 2023; Yu et al., 2022; NVIDIA, 2024b; HuggingFace, 2024; Qiu et al., 2024a; Li et al., 2023; Sheng et al., 2023b; Wu et al., 2023b) largely focused on serving interactive requests, such as chatbots, with tight latency SLO requirements. However, the recent explosive growth of LLM applications, has generated a need to support batch LLM queries with SLO values ranging from minutes to hours for tasks such as data wrangling (Narayan et al., 2022), document processing (Jin et al., 2024), and model fine-tuning (Sudalairaj et al., 2024).
Given the broader range of SLO requirements and the use of multiple models, maintaining request queues is beneficial. Conceptually, when resources are limited, requests with relaxed SLOs can be kept at the back of the request queue and do not need to be executed immediately, while requests with tight SLOs can be kept ahead in the queue to ensure immediate execution and prevent head-of-line (HOL) blocking. Hence, queue ordering becomes an important decision-making problem for SLO-oriented LLM serving.
Previous work in SLO-oriented serving has largely focused on traditional DNN serving workloads such as CNNs and RNNs (He et al., 2016; Salehinejad et al., 2017), where the queue ordering decisions are made by systems such as Clockwork (Gujarati et al., 2020), INFaaS (Romero et al., 2021), and SHEPHERD (Zhang et al., 2023b). Such systems leverage the deterministic execution times of DNN workloads to estimate queuing times and enable optimal ordering decisions that would maintain request SLOs. However, such systems cannot be easily applied to LLM serving workloads because the execution time per request is non-deterministic as the number of output tokens are unknown apriori. Using these system’s assumption of fixed size batches with deterministic execution times for scheduling decisions leads to sub-optimal scenarios, described below.
Figure 1. Previously proposed SLO-oriented serving systems overestimate queue waiting time leading to suboptimal resource usage. (Left) Estimated waiting time when requests are run with Llama-70B on A100 GPUs with vLLM. (Right) Number of GPUs required to maintain 20s time-to-first-token (TTFT) SLO with previous systems vs QLM in single and multi-model scenarios.
[…]
Previously proposed SLO-oriented serving systems overestimate queue waiting time leading to suboptimal resource usage. (Left) Estimated waiting time when requests are run with Llama-70B on A100 GPUs with vLLM. (Right) Number of GPUs required to maintain 20s time-to-first-token (TTFT) SLO with previous systems vs QLM in single and multi-model scenarios.
Figure 2. QLM uses request groups and LLM Serving Operations (LSOs) such as request eviction to minimize resource requirement. Previously proposed systems would use four vLLM instances (compared to two for QLM) due to limitations described in Figure 1.
[…]
QLM uses request groups and LLM Serving Operations (LSOs) such as request eviction to minimize resource requirement. Previously proposed systems would use four vLLM instances (compared to two for QLM) due to limitations described in Figure 1.
First, previously proposed serving systems are unable to effectively multiplex interactive and batch requests on the same serving instance. We find that the queue waiting time in LLM serving is lower than estimate used by the scheduler in such systems as shown in Figure 1 (left). As these systems generate a higher than actual estimate of the total queuing time, they allocate batch and interactive requests on separate serving instances, even in scenarios where allocation on the same instance would suffice to meet SLOs.
Second, previously proposed systems are unable to effectively multiplex different models on the same serving instance. Similar to the first scenario, systems such as Clockwork (Gujarati et al., 2020) overestimate the time required to serve a single model and prefer to allocate independent instances for each model to preserve SLOs. Others such as SHEPHERD (Zhang et al., 2023b), altogether forgo serving multiple models on the same instance due to the high cost of model swapping relative to inference times.
Our Work. To address these limitations, we propose QLM, a queue management system for LLM serving that maximizes SLO attainment while maintaining high throughput. At the core of QLM is the Request Waiting Time (RWT) Estimator (described in Section 6) that estimates the waiting times for requests in the request queue. We observe that as the queue size grows larger, statistical effects of continuous batching (Yu et al., 2022) in LLM serving allows us to create a tighter bound on waiting time. Additionally, we empirically validate and provide a proof for applicability of the RWT estimator. QLM leverages this improved estimation to allow for better utilization and decrease resource cost by making closer to optimal request queue reordering decisions.
To order requests in the queue, QLM, similar to SHEPHERD, uses the abstraction of request groups (described in Section 4). Each request group is a collection of requests that have relatively homogeneous performance requirements (e.g. similar SLOs and model). QLM places these request groups in virtual queues that determine the order in which requests are consumed by the LLM serving instances. In the LLM serving context, in addition to the benefits of scalability and predictability found in DNN serving, request groups are a useful abstraction for model swapping. As decision making is made at the per-request group level, the total amount of swapping is minimized and throughput is improved as depicted in Figure 2 (Scenario 2).
Finally, QLM uses a Global Scheduler (described in Section 7) that utilizes the waiting time estimates of request groups to create an optimal ordering and assignment of request groups onto the virtual queues to maximize SLO attainment. The virtual queue ordering is used by four basic LLM Serving Operations (LSOs) to manage the request queue (see section 5 for details): (1) Request Pullingfrom the global waiting queue into the running batch in the GPU, (2) Request Evictionfrom the running batch back into the waiting queue, (3) Model Swappingfrom CPU to GPU memory, and (4) Load Balancingacross multiple LLM model instances. For example, request eviction allows serving batch and interactive requests on the same LLM serving instance and prevents HOL blocking for interactive requests as shown in Figure 2 (Scenario 1).
@@ 1. Introduction
Results. We demonstrate QLM on vLLM (Kwon et al., 2023) as the backend LLM-serving system. We evaluate QLM on three popular LLMs of varying sizes (i.e., Mistral-7B (Jiang et al., 2023), Vicuna-13B (Chiang et al., 2023), and Llama-70B (Touvron et al., 2023)) on GPU clusters with NVIDIA A10 and A100 GPUs. We adopt workloads from a real-world LLM dataset: ShareGPT (team, 2024) using setups derived from our production requirements. Our experiments demonstrate the following major improvements with QLM:
@@
    SLO Attainment: Depending on the arrival rate, QLM achieves 40–90% higher SLO attainment compared to the vanilla vLLM serving system and 50–90% higher SLO attainment compared to traditional ML serving systems like SHEPHERD.
@@
    Request Throughput: QLM improves the request throughput in a multi-model serving system by 400% on average and in a single-model serving system by 20% on average compared to other LLM serving systems.
@@
    LSO Ablation Study: QLM demonstrates that all LSOs contribute to SLO attainment and throughput improvement. Notably, we find that model swapping improves throughput by 300% in multi-model serving, and request eviction improves SLO attainment by 80% in single-model serving.
QLM has been merged into an internal production LLM routing service.
2. Background
2.1. LLM Inference
@@
Continuous Batching. During LLM inference, the decoding stage is memory-bound, as loading model weights from memory takes longer than computation. Therefore, state-of-the-art LLM serving systems like vLLM (Kwon et al., 2023), Orca (Yu et al., 2022), Tensor-RT (NVIDIA, 2024b) and TGI (HuggingFace, 2024) employ continuous batching with iterative scheduling to enable dynamic addition of requests to a batch as soon as others have finished generation.
@@
PagedAttention. Static allocation of the KV cache can result in significant memory waste as the KV cache grows dynamically during the decoding stage. PagedAttention (Kwon et al., 2023) introduces the idea of managing the KV cache, like OS memory, via pages and enabling dynamic allocation. Such dynamic allocation prevents fragmentation and enables nearly 100% utilization of GPU memory and furthers throughput improvement when combined with continuous batching (Kwon et al., 2023).
Figure 3. Requests have predictable waiting times in a continuous batching system.
[…]
Requests have predictable waiting times in a continuous batching system.
Figure 4. Forced request eviction leads to reduction in head-of-line (HOL) blocking time.
[…]
Forced request eviction leads to reduction in head-of-line (HOL) blocking time.
Figure 5. Model swapping and request pulling can jointly decrease queue drain time.
[…]
Model swapping and request pulling can jointly decrease queue drain time.
2.2. LLM Serving Systems
@@
User-facing applications such as chatbots, log processing, and recommenders have specific latency SLO requirements (Qiu et al., 2020; Jyothi et al., 2016; Qiu et al., 2024b; Chen et al., 2019). When interacting with LLMs, each application generates requests that consist of the input prompts and associated metadata (e.g., model type and SLO value) to the LLM serving system. For example, chatbots require requests to complete by a deadline (e.g., p99 time to first token (TTFT) ¡ 20s (Gnewuch et al., 2022)), and batch jobs like document processing have a more relaxed SLO in the order of minutes to hours. Requests may also need to be served by multiple fine-tuned models specialized for various tasks. For example, Code Llama (Roziere et al., 2023) is fine-tuned for coding assistance, and Llama-chat (Touvron et al., 2023) is fine-tuned for chatbots. Maintaining a standalone LLM serving system for each of these models or SLO types can be expensive, and they often have to be multiplexed together to share the same serving system (Choi et al., 2022; Li et al., 2023; LeMay et al., 2020; Sheng et al., 2023a). Multiplexing batch and interactive requests across various models with limited resources leads to formation of request queues, and managing these queues is critical to comply with SLO requirements.
2.3. QLM Definitions
@@
SLO: Each request arrives with a service-level objective (SLO) value that enforces the time to first token (TTFT) for the request. While QLM primarily focuses on TTFT, it can be paired with another system such as Andes (Liu et al., 2024) to also maintain inter-token latency (ITL).
@@
LLM Serving Instance: An LLM serving system [^1] The difference between an LLM serving system and an LLM serving instance is similar to JAVA classes and objects. For example, vLLM is an LLM serving system while vLLM with a loaded model like Llama 70B is an instance of the LLM serving system. is capable of hosting LLM models by providing the necessary infrastructure and resources to load the models into memory and respond to requests. QLM is compatible with existing LLM serving systems such as vLLM (Kwon et al., 2023) and TGI (HuggingFace, 2024). An LLM serving instance is composed of the LLM serving system and an LLM model that is being served.
2.4. Motivation and Characterization
To enable SLO-oriented LLM serving, it is critical to understand (1) the impact of LLM autoregressive patterns on the request waiting time, (2) the performance implications of queuing batch and interactive requests, and (3) the performance implications of queuing multi-model batch requests.
@@ 2.4. Motivation and Characterization
We characterize these scenarios with a state-of-the-art LLM serving system, vLLM (Kwon et al., 2023), to motivate the design of QLM. We use ShareGPT (team, 2024) traces to evaluate the system. We present three key insights below.
Insight #1: Waiting times in long request queues can be accurately estimated analytically. Recall from Figure 1 that incorrectly estimating waiting times can lead to resource waste. Therefore, we attempt to build an accurate queue waiting time estimator.
While the completion time for individual requests in LLM inference can vary widely, the average waiting time for a request in a long queue is predictable. Due to the statistical averaging effects resulting from a large number of requests (as detailed in Appendix A.1), the waiting time can be estimated by dividing the total number of output tokens for the pending requests by the token generation throughput. The total number of output tokens adheres to a Normal distribution, in accordance with the Central Limit Theorem, since individual requests are independent of each other
Figure 5 illustrates this linear relationship between waiting time and queue position when serving requests for three varying-sized LLMs on NVIDIA A100 GPUs. Additionally, we find that the estimator is highly accurate with a coefficient of determination ($R^{2}$) of 0.99 (out of 1.0).
Insight #2: HOL blocking times due to continuous batching can be in the order of tens of seconds. The straightforward way to prioritize interactive requests over batch requests on the same LLM serving instance is by placing them at the front of the waiting queue. However, request placement in the waiting queue may not be sufficient for immediate execution due to the lack of available GPU memory, causing head-of-line (HOL) blocking. Therefore, in such scenarios, evicting batch requests from the GPU is required. To minimize the cost of eviction, we can preserve the KV cache of batch requests, allowing execution to resume from the last decoding iteration.
Figure 5 illustrates the HOL blocking time when run with a mixed workload comprising interactive and batch requests. In the absence of request eviction, the HOL blocking time can be in the order of several seconds, which can lead to violation of latency SLOs for interactive requests. However, request eviction significantly reduces the waiting time because interactive requests only need to wait for a single decoding iteration before they can be scheduled, resulting in a 100–1000$\times$ reduction in waiting time.
@@
Insight #3: Policies such as Earliest Deadline First (EDF) are insufficient to eliminate HOL blocking from model swapping. The optimal request pulling strategy to maximize the number of requests that satisfy SLOs is Earliest Deadline First (EDF) scheduling. However, this assumes that the model swapping cost is negligible. Frequent model swaps can happen (similar to thrashing) if multiple models are served to time share the same GPU devices, leading to SLO violations due to longer completion times to drain the queue and a drop in throughput. For example, consider the case illustrated in Figure 5. Requests with varying SLOs arrive in the queue, and they are placed by an EDF policy, causing multiple model swaps and substantially higher time to drain the entire request queue. Specifically, we find that across models and GPUs, the time required to serve all requests in the queue (i.e., the queue drain time) is substantially higher for the EDF policy compared to an Oracle policy that groups requests from the same model together to prevent the overhead of repetitive model swaps.
3. QLM Design Overview
Figure 6. Overview of QLM.
[…]
Overview of QLM.
QLM aims to maximize latency SLO satisfaction in LLM serving workloads. To do so, QLM manages a global request queue and orchestrates multiple LLM Serving Operations (LSOs) to reorder and drain the global queue.
3.1. Lifecycle of a Request in QLM